Greenhab Report – January 11th

GreenHab Report

Mark Gee


Environmental control:


Shade cloth on)

Working Hour: 06:30PM Inside temp at working hour: 17 C Outside temp during working hours: 3 C Inside temperature high: 33 C Inside temperature low: 16 C Inside humidity: 39 %RH

Inside humidity high: 41 %RH Inside humidity low: 16 %RH

Hours of supplemental light:

For the crops 05:00 to 11:59 PM

Changes to crops: The tomatoes, cucumbers, beans, herbs, peppers, and melons keep growing. The microgreens have taken off and will be ready to harvest tomorrow. Of the seeds planted, the arugula, mustard, and various lettuce varieties have sprouted.

Daily water usage for crops: 15 gallons

Time(s) of watering for crops: 11:00AM, 06:30PM

Research observations: The microgreens have recovered from the early damage and are sprouting magnificently. Tomorrow they will be harvested.

Changes to research plants: None.

Aquaponics: Not in use.

Narrative: Today I watered the plants as usual and inventoried the tools, fertilizer, pots, and supplies.

The grow lights turned on at 02:00 pm. The analog timer has been gaining time steadily, as do the clocks on the electronic appliances in the Hab. These clocks probably use the frequency of the alternating current power supply to keep time as is common in most appliances. The standard frequency for an electrical grid in the US is 60 Hz, and our crew engineer observed that the generator is supplying power at 60.18 Hz. This increased frequency would cause the clocks to gain time at approximately the rate we are observing. See Sol 1 summary for a further explanation.

To compensate for this gain, the light timer needs to be reset every day.

Support/supplies needed: Scale with a capacity of 2 kg. The analog scale in the Science Dome has a capacity of 610g. Needed for experiments tomorrow.

EVA Report – January 11th

Author: Max Fagin

Purpose of EVA: The last testing of the Yagi-Uda antenna based navigation experiment for navigating back to the hab. Destination change the night before to Skyline ridge allowed photographic reconnaissance in preparation for tomorrow’s geology EVA to the same site.

Destination: Plains to the east of skyline ridge

UDM27 Coordinates: 516000 E, 4251000 N

Participants: Max Fagin (Commander), Kshitij Mall (Executive Officer)

Narrative: This EVA was the final demonstration of the radio beacon navigation system, with the goal of demonstrating its effectiveness when line of sight was obstructed for the duration of the walk. Kshitij served as the lost astronaut, and had the hood in place over his helmet that limited his field of view to 5m from his feet. Max monitored the route and tracked the progress via gps to ensure the route was safe. We left the airlock at 10:36 and walked north along a foot path to the top of the ridge that rises over the hab to our west. After cresting the ridge, we walked due west along Sagan Rd before swinging south 50m to await the first beacon transmission. When the beacon was received, we started back, Kshitij leading the way, and Max monitoring. When Kshitij would stop to check the bearing on the hab, Max tested an experimental chlorophyll detector that Chris McKay and Matteo Borri had designed for us for use on the local flora. A loss of the beacon for ~15 minutes led to a complete turnaround, but it was corrected when the beacon was picked up again. We crested the ridge 45 minutes later at the ridge’s closest point to the hab, exactly where we should have been. So close that Kshitij gave an audible gasp when Max removed the hood and saw the habitat was so close. We deactivated the beacon, and walked north to find the route back down to the hab. Airlock ingress was at 12:37 for a total EVA time of 2:01.

Note: One of the hab’s radios has been damaged beyond repair. I was carrying it in my front pocket as a backup radio, and when I lifted my leg while climbing the ridge, the antenna was pinched against my front and snapped off. Shannon: There are still more than enough radios at the hab to provide spares, so I don’t know if you will want to replace it or not, but either way, I will note this on the checkout sheet as damage incurred. My apologies for the damage.

Sol Summary – January 11th

Crew 186 Sol 11 Summary Report 11JAN2018

Sol 11

Summary Title: The first of the lasts

Author’s name: Max Fagin

Mission Status: Preparing for a grand finale EVA to a geology site we have saved especially for our last EVA.

Sol Activity Summary: The sun came back out today, much to our appreciation (and the appreciation of our solar panels). Any problem can be fixed in a spacecraft as long as you still have power and comms, and it has been a great weight off our shoulders to only view cloudy days as a mild annoyance, rather than a critical power threat as they might be on a pure solar mission.

The clear weather allowed for our last navigation EVA to take place as scheduled. We now have four complete demonstrations of an astronaut using nothing but the radio beacon to safely return to the hab, two where the astronaut had no experience with the system, and an impaired range of vision. I’m very pleased with Justin for producing such a practical aid to navigation in places where no gps is available (such as Mars … for now at least). It’s an early back to school for some of us, as Melanie has begun taking us on short engineering EVAs to test the efficacy of the VR training module she has developed for teaching EVA tasks. Around the hab, we continue to check off the last remaining items on our individual projects so we can face handoff with a clear schedule.

On a personal note, today was the first day I really saw GreenHab as more than just a greenhouse. I remember contributing to the Kickstarter campaign to rebuild Greenhab after the fire, and it has made me immeasurably proud to see it continue to thrive and grow under Mark’s skillful control while we are here. Stepping into the tropical heat and humidity of GreenHab today after a cold and dry EVA felt as relaxing as stepping into a hot tub. I sat down amidst the plants to appreciate the magnificent oasis thriving in this red/brown landscape, and the environment was so placid that I was not at all surprised when I woke up 15 minutes later to find I had drifted to sleep in my chair. Greenhouse, oasis, zen garden, GreenHab is much more now than just a place where plants grow.

Look Ahead Plan: Our last EVA will take place tomorrow, with one final geological survey to the highest elevation site on our list at the base of Skyline Ridge. Harvesting of our carefully grown GreenHab bounty will also take place tomorrow, we are planning to exit sim at noon on Saturday 01/13.

Anomalies in work: None

Weather: 6C, Wind SSE 8km/h, Fair, Humidity 52%

Crew Physical Status: Healthy

EVA: Preliminary photography of the Plane of Dirt / Skyline Ridge region for the last geology EVA, then walk back to the hab to test the radio navigation system.

Reports to be filed: Sol Summary, Operations Report, GreenHab Report, Journalist Report, EVA Report, EVA Request

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site